# Self-Evolved Reward Learning for LLMs

### v1

@@ 3.1 Overview
Furthermore, by allowing the LLM to generate two answers for each question, the RM is provided with paired examples that are key to both learning statuses. This setup enriches the training data with diverse answer qualities, enabling the RM to improve its discrimination and comparative abilities. Once the RM becomes proficient at handling both tasks, it is well-equipped to guide the LLM during reinforcement learning.
@@
$\mathcal{S}=\begin{cases}\text{Status}_{1},&\text{if }(p_{i}^{1}>\tau_{\text{high}}\text{ and }p_{i}^{2}<\tau_{\text{low}})\text{ or }(p_{i}^{1}<\tau_{\text{low}}\text{ and }p_{i}^{2}>\tau_{\text{high}}),\\ \text{Status}_{2},&\text{if }\Delta_{p}\geq\tau_{\Delta},\\ \text{Stop},&\text{otherwise}.\end{cases}$ (3)
In equation 3, Status 1 means the model can confidently distinguish between two answers when one has a high probability of being good ($p_{i}^{1}>\tau_{\text{high}}$) and the other a low probability ($p_{i}^{2}<\tau_{\text{low}}$), or vice versa. Status 2 means the model refines its judgment when the probability difference $\Delta_{p}$ exceeds $\tau_{\Delta}$, useful when both answers are either good or bad but differ in relative quality. Stop: no further self-training if neither condition is met, signaling the model has reached its capacity on the current data. This approach allows the model to first learn clear distinctions between good and bad answers (Status 1) before focusing on more subtle differences (Status 2). When improvements from Status 1 are marginal, we shift to Status 2 to refine the model’s ability to handle nuanced cases.
Why capturing different statuses? The different statuses reflect the complexity of the relationships between $A_{i}^{1}$ and $A_{i}^{2}$. The reward model needs to manage both clear distinctions and subtler comparisons. By identifying the model’s learning status $\mathcal{S}$, we ensure that the model trains on the most relevant data for its current focus—strengthening its ability to handle clear distinctions or refining its comparative abilities. This staged approach maximizes the model’s learning efficiency, first focusing on easier tasks (Status 1) and then moving on to more complex judgments (Status 2). Experiments validate that alternating between these statuses leads to improved performance overall.
@@
Here, $D_{\text{unlabeled}}$ refers to the unlabeled data. In Status 1, the filter selects high-confidence data where $(RM(Q_{j},A_{j}^{1})>\tau_{\text{high}}\text{ and }RM(Q_{j},A_{j}^{2})<\tau_{\text{low}})$ or $(RM(Q_{j},A_{j}^{1})<\tau_{\text{low}}\text{ and }RM(Q_{j},A_{j}^{2})>\tau_{\text{high}})$, ensuring the model trains on reliable examples. In Status 2, the filter selects pairs where the reward difference $|RM(Q_{j},A_{j}^{1})-RM(Q_{j},A_{j}^{2})|$ exceeds a threshold $\delta$, focusing on refining comparative judgments. After filtering, the model is retrained using pairwise loss, allowing the model to compare answers relatively rather than relying on absolute labels. which consistently improves performance by focusing on relative comparisons rather than absolute classifications. The pairwise loss function is:
@@
where $\Delta$ is the desired margin between reward scores. This iterative process, i.e., filtering data and retraining with pairwise loss, enables the RM to progressively refine its judgment until it converges.
@@ 4.1 Reward Modeling Results
Table 1: The results of reward modeling on the HH-RLHF, Ultrafeedback, Summarize, and Stackoverflow. Loop 0 denotes the RM trained with 15% of the human data. Our method represents the results of iterative evolution based on the Loop 0 model. Full dataset denotes the results of the RM trained with the entire set of human-annotated data.
@@ 4.1.1 Main Findings
SER consistently and effectively enhances model performance. As shown in Table 1, compared to the baseline that uses only 15% of the data, SER improves the model’s performance by incorporating self-labeled data for training. Through multiple iterations, the model achieves significant performance gains, resulting in an average 7.88% increase in accuracy. Furthermore, we find that as the model’s parameter size increases, the foundational capability strengthens, and the potential for SER’s self-improvement further enhances. Larger parameter models can achieve higher self-improvement benefits. Mistral 7B attains an average performance improvement of 7.74% during iterations, Llama 8B achieves an average improvement of 8.94%, and Llama 13B sees an average improvement of 6.76%. Across four datasets, Llama 8B surpasses Mistral 7B.
@@
SER can approach or even exceed the performance of full-scale human-labeled data. We compare our method with using the full human-labeled data. The results demonstrate that SER can achieve performance close to that of using the complete human-labeled dataset, with an average performance difference of 0.3%. For Mistral 7B on the HH-RLHF dataset, SER exceeds the baseline by 0.13%, and on the Summarize dataset, it surpasses the baseline by 1.93%. For LLaMA 8B on the UltraFeedback dataset, it achieves a performance advantage of 0.54% over the baseline. Additionally, a trend is observed where the disparity between the SER method and the full human-labeled data increases with the model parameter size. Specifically, Mistral 7B exceeds the baseline by an average of 0.12%, LLaMA 8B by 0.06%, while for LLaMA 13B, the baseline surpasses SER by an average of 1.07%. This indicates that as the model parameter size increases, its capacity to accommodate more labeled data also increases, thereby enhancing performance. This is a positive trend, suggesting that we can further enhance the performance ceiling of models by obtaining more labeled data through SER rather than manual labeling.
@@ 4.2 PPO Results
To validate the effectiveness of SER, we use the previously mentioned RM to guide PPO training, thereby optimizing the LLM. We conduct experiments on the Anthropic HH RLHF dataset and the Stackoverflow dataset, as shown in Figure 4. In the hh-rlhf dataset, all self RLHF models exceed the SFT baseline in terms of win rate, indicating that the self RLHF approach enhances the capabilities of the LLM. Compared to RMs trained with the full human-annotated data, the win rate in PPO experiments demonstrates a consistent trend with the performance of the RMs. For Mistral 7B, the accuracy of the self RLHF RM surpasses that of the RM trained with the full dataset, and in PPO experiments, the win rate also slightly exceeds that of the full model. Additionally, to verify the generalizability of our method, we conduct the same experiment on the StackOverflow dataset. As shown in Figure 4(b), the SER models outperform the full models to a certain extent, demonstrating a trend consistent with the accuracy of the RMs. In summary, our main findings are as follows:
@@ Appendix A THEORETICAL ANALYSIS
In this section, we provide a detailed theoretical analysis to support the effectiveness of our Self-Evolved Reward Learning (SER) method. The analysis focuses on two key aspects:
@@
    Convergence of the Reward Model during Self-Training: We analyze conditions under which the RM improves through iterative self-training based on predicted probabilities.
@@
    Convergence Properties of PPO with a Learned Reward Model: We discuss the theoretical properties of using PPO with an RM that has been trained using the self-evolved process.
@@ A.1 Convergence of the Reward Model during Self-Training
In our SER framework, the RM iteratively improves itself by filtering self-labeled data based on predicted probabilities and retraining using pairwise loss. The theoretical analysis focuses on the convergence of the RM’s predicted probabilities over time.
A.1.1 Assumptions
Assumption 1 (Initial Quality of Predicted Probabilities).
The initial RM $R_{\theta}^{(0)}$ provides predictions where, for a pair of answers $(A_{i}^{1},A_{i}^{2})$, the probability difference $\Delta_{p}^{(0)}=|p_{i}^{1}-p_{i}^{2}|$ is meaningful, i.e., the probability difference reflects the relative quality of answers with reasonable accuracy.
Assumption 2 (High-Confidence Data Filtering).
During self-training, only pairs of answers $(A_{i}^{1},A_{i}^{2})$ where the predicted probability difference exceeds a confidence threshold $\tau_{p}$ (e.g., $|p_{i}^{1}-p_{i}^{2}|>\tau_{p}$) are used for further training. This ensures that the model uses reliable self-labeled data.
A.1.2 Theoretical Result
@@
Under Assumptions 1 and 2, and with appropriate choices of confidence threshold $\tau_{p}$ and pairwise comparison margin $\delta$, the self-training process of the RM converges, and the predicted probabilities become increasingly reliable with each iteration, i.e., the RM improves over time.
A.1.3 Proof Sketch
The key steps of the proof are as follows:
  1. 1.
    Initial Quality of Predictions: From Assumption 1, the RM begins with predictions that, while not perfect, provide meaningful probability differences between answer pairs.
  2. 2.
    High-Confidence Data Selection: By filtering the data based on the confidence threshold $\tau_{p}$ (Assumption 2), we ensure that only reliable, high-quality data points are used for retraining. This avoids the risk of propagating errors from uncertain self-labeled data.
  3. 3.
    Iterative Improvement: Each iteration of the self-training process uses pairwise loss to refine the RM’s ability to distinguish between answer pairs, resulting in improved predicted probabilities with each training step.
  4. 4.
    Convergence: As the self-training process continues, the RM’s predicted probabilities become more accurate, converging to a stable state where further self-training leads to diminishing returns in performance improvement.
A.1.4 Discussion
Theorem 1 shows that, under reasonable conditions, the self-training process for the RM leads to improved performance over time. The key is to carefully filter the data based on predicted probabilities to ensure that only reliable self-labeled examples are used for training. This filtering mechanism prevents error propagation and ensures that the RM improves iteratively.
@@ A.2 Convergence Properties of PPO with a Learned Reward Model
Proximal Policy Optimization (PPO) is used to train the LLM with reward signals generated by the self-evolved RM. The theoretical challenge lies in ensuring that PPO can converge effectively when using an RM that has been trained through self-labeling, which may introduce some inaccuracies in the reward estimates.
A.2.1 Assumptions
Assumption 3 (Bounded Reward Estimation Error).
The RM provides reward signals where the error relative to the true reward function $R^{*}(Q,A)$ is bounded, i.e., $|RM(Q,A)-R^{*}(Q,A)|\leq\epsilon_{r}$, where $\epsilon_{r}$ is small.
A.2.2 Theoretical Result
Theorem 2 (Policy Improvement with Approximate Rewards).
Under Assumption 3 and using PPO with appropriately chosen hyperparameters, the LLM’s policy $\pi_{\phi}$ converges to a near-optimal policy with respect to the true reward function $R^{*}$, with the error bounded by $\epsilon_{r}$.
A.2.3 Proof Sketch
The proof proceeds as follows:
  1. 1.
    Consistency of Policy Gradient: The policy gradient computed using the learned RM is close to the true policy gradient computed using $R^{*}$, with the error bounded by $\epsilon_{r}$.
  2. 2.
    Effect of Reward Estimation Error: The bias introduced by the reward estimation error affects the gradient update but remains small, as long as $\epsilon_{r}$ is small.
  3. 3.
    PPO Stability: PPO’s clipping mechanism prevents large updates, ensuring that policy updates remain stable even in the presence of small reward estimation errors.
  4. 4.
    Convergence: By iteratively applying PPO updates, the policy converges to a near-optimal policy relative to the true reward function, with the error in the final policy bounded by $\epsilon_{r}$.
A.2.4 Discussion
Theorem 2 shows that PPO remains effective even when using a learned RM, provided the reward estimation errors are small. Our self-evolved RM minimizes $\epsilon_{r}$ over time by iteratively refining its predicted probabilities, ensuring that the reward signals used for policy optimization become more accurate. As a result, PPO can still guide the LLM toward a near-optimal policy.
Appendix B Algorithm Summary

### v2

@@ 3.1 Overview
Furthermore, by allowing the RM to judge two answers for each question, the RM is provided with paired examples that are key to both learning statuses, enabling the RM to improve its discrimination and comparative abilities. Once the RM becomes proficient at handling both tasks, it is well-equipped to guide the LLM during reinforcement learning.
@@
$\mathcal{S}=\begin{cases}\text{Status}_{1},&\text{if }(p_{i}^{1}>\tau_{\text{high}}\text{ and }p_{i}^{2}<\tau_{\text{low}})\text{ or }(p_{i}^{1}<\tau_{\text{low}}\text{ and }p_{i}^{2}>\tau_{\text{high}}),\\ \text{Status}_{2},&\text{else if }\Delta_{i}\geq\tau_{\Delta},\\ \text{Stop},&\text{otherwise}.\end{cases}$ (3)
To determine the current status, we use the reward model (RM) trained in the current iteration to predict on the unlabeled data. Both Status 1 and Status 2 require a sufficient number of predictions meeting specific criteria to ensure a statistically meaningful assessment. (In this paper, we selected $\tau_{\text{high}}=0.55$, $\tau_{\text{low}}=0.45$, and $\tau_{\Delta}$ = 0.3 as they provided the most consistent improvements in the RM’s ability)
  • Status 1 (Easier Task): This status evaluates whether the RM can effectively distinguish between positive (good) and negative (bad) samples. The evaluation is based on the predicted probabilities $p_{j}^{k}$ for each answer $A_{j}^{k}$: If $p_{j}^{k}>\tau_{\text{high}}$, the RM is confident that the answer is positive. If $p_{j}^{k}<\tau_{\text{low}}$, the RM is confident that the answer is negative. A sufficient number of high-confidence predictions (e.g., 600 in the HH dataset) indicates that the RM is proficient in distinguishing positive and negative samples, thereby satisfying the criteria for Status 1.
  • Status 2 (Harder Task): This status assesses the RM’s ability to discern subtle differences between answers of similar quality (e.g., both good or both bad). It requires the RM to evaluate paired answers to the same question and compute the absolute difference between their predicted probabilities:
    $\left|p_{j}^{1}-p_{j}^{2}\right|>\tau_{\Delta}$
    If a sufficient number of paired predictions meet this threshold, it indicates that the RM can amplify distinctions between similar-quality answers. This task is more challenging than Status 1 because it requires the RM to recognize and quantify nuanced differences. Similar to Status 1, this determination requires a sufficient number of predictions on the unlabeled dataset (e.g., 600 predictions in the HH dataset).
We check the statuses in order—first Status 1 and then Status 2—because Status 1 represents a foundational capability that is necessary before tackling the more complex task in Status 2. Status 1 is the easier task, focusing on broad distinctions, while Status 2 is the harder task, requiring finer-grained analysis. If the RM does not meet the criteria for Status 1 (i.e., few or no samples satisfy the thresholds $\tau_{\text{high}}$ and $\tau_{\text{low}}$), we then check for Status 2. If the RM also fails to meet the criteria for Status 2, we interpret this as the model reaching its convergence point, and we halt further training of the RM.
@@
Here, $D_{\text{unlabeled}}$ refers to the unlabeled data. In Status 1, the filter selects high-confidence data where $(RM(Q_{j},A_{j}^{1})>\tau_{\text{high}}\text{ and }RM(Q_{j},A_{j}^{2})<\tau_{\text{low}})$ or $(RM(Q_{j},A_{j}^{1})<\tau_{\text{low}}\text{ and }RM(Q_{j},A_{j}^{2})>\tau_{\text{high}})$, ensuring the model trains on reliable examples. In Status 2, the filter selects pairs where the reward difference $|RM(Q_{j},A_{j}^{1})-RM(Q_{j},A_{j}^{2})|$ exceeds a threshold $\delta$, focusing on refining comparative judgments.
After filtering, the model is retrained using pairwise loss, allowing the model to compare answers relatively rather than relying on absolute labels. which consistently improves performance by focusing on relative comparisons rather than absolute classifications. The pairwise loss function is:
@@
where $\Delta$ is the desired margin between reward scores, and ${D_{\text{filtered}}}={D^{n}_{filtered}}+{D^{n-1}_{filtered}}$, where $n$ denote the number of iterations of the loop. The training data for the current loop consists of the data filtered using Equation 4, in addition to the training data from the previous loops. This iterative process, i.e., filtering data and retraining with pairwise loss, enables the RM to progressively refine its judgment until it converges.
@@ 4.1 Reward Modeling Results
Table 1: The results of reward modeling on the HH-RLHF, Ultrafeedback, Summarize, and Stackoverflow. Loop 0 denotes the RM trained with 15% of the human data. SER represents the results of iterative evolution based on the Loop 0 model. Full dataset denotes the results of the RM trained with the entire set of human-annotated data.
@@ 4.1.1 Main Findings
SER consistently and effectively enhances model performance. As shown in Table 1, compared to the baseline that uses only 15% of the data, SER improves the model’s performance by incorporating self-labeled data for training. Through multiple iterations, the model achieves significant performance gains, resulting in an average 7.88% increase in accuracy. Furthermore, we find that as the model’s parameter size increases, the foundational capability strengthens, and the potential for SER’s self-improvement further enhances. Larger parameter models typically achieve higher performance after undergoing self-improvement. In most experimental settings, the performance of the LLama 13B model surpasses that of the other two smaller parameter models.
@@
SER can approach or even exceed the performance of full-scale human-labeled data. We compare our method with using the full human-labeled data. The results demonstrate that SER can achieve performance close to that of using the complete human-labeled dataset, with an average performance difference of 0.3%. For Mistral 7B on the HH-RLHF dataset, SER exceeds the baseline by 0.13%, and on the Summarize dataset, it surpasses the baseline by 1.93%. For LLaMA 8B on the UltraFeedback dataset, it achieves a performance advantage of 0.54% over the baseline. A potential trend is observed where the difference between the SER method and the full human-labeled data increases with model size. Specifically, the average difference for Mistral 7B is +0.12%, for LLaMA 8B is +0.06%, and for LLaMA 13B is -1.07%. This suggests that larger models better utilize labeled data, enhancing performance. This trend highlights the potential of SER to further elevate model performance by scaling labeled data through self-labeling rather than manual annotation.
@@ 4.2 PPO Results
To validate the effectiveness of SER, we use the previously mentioned RM to guide PPO training, thereby optimizing the LLM. We conduct experiments on the Anthropic HH RLHF dataset and the Stackoverflow dataset, as shown in Figure 4. In the hh-rlhf dataset, all SER models exceed the SFT baseline in terms of win rate, indicating that the SER approach enhances the capabilities of the LLM. Compared to RMs trained with the full human-annotated data, the win rate in PPO experiments demonstrates a consistent trend with the performance of the RMs. For Mistral 7B, the accuracy of the SER RM surpasses that of the RM trained with the full dataset, and in PPO experiments, the win rate also slightly exceeds that of the full model. Additionally, to verify the generalizability of our method, we conduct the same experiment on the StackOverflow dataset. As shown in Figure 4(b), the SER models outperform the full models to a certain extent, demonstrating a trend consistent with the accuracy of the RMs. In summary, our main findings are as follows:
@@ Appendix A Theoretical Analysis
In this section, we provide a rigorous theoretical analysis to support the effectiveness and convergence of our SER method. We focus on two main aspects:
@@
    Convergence of the Reward Model during Self-Training: We provide formal conditions under which the reward model converges to a good solution through self-training.
@@
    Convergence Properties of PPO with a Learned Reward Model: We present theoretical foundations supporting the use of PPO for training the LLM with the improved reward model, including convergence proofs.
@@ A.1 Convergence of the Reward Model during Self-Training
Self-training involves using a model’s own predictions to generate additional training data. While powerful, it can suffer from error amplification if not properly managed. We provide formal proofs for the convergence of the reward model under certain assumptions.
Definitions
Let $R_{\theta}^{(t)}$ denote the reward model at iteration $t$. Let $\mathcal{D}_{\text{filtered}}^{(t)}$ be the filtered dataset used for retraining at iteration $t$.
Assumptions
Assumption 1 (Initial Model Accuracy).
The initial reward model $R_{\theta}^{(0)}$ has an expected accuracy greater than random guessing:
$\mathbb{P}_{(Q,A,y)}\left[R_{\theta}^{(0)}(Q,A)=y\right]=\text{Acc}^{(0)}>0.5,$ (7)
where $y\in\{0,1\}$ denotes the true label (bad or good answer).
Assumption 2 (High-Confidence Prediction Reliability).
For data points where the reward model’s prediction confidence exceeds thresholds $\tau_{p}$ and $1-\tau_{p}$, the prediction accuracy is at least $\alpha$, with $\alpha>0.5$:
$\mathbb{P}\left[R_{\theta}^{(t)}(Q,A)=y\mid|p^{(t)}(Q,A)-0.5|\geq\delta_{p}\right]\geq\alpha,$ (8)
where $p^{(t)}(Q,A)$ is the predicted probability, and $\delta_{p}=\tau_{p}-0.5$, $\tau_{p}$ equala to $\tau_{high}$ and $1-\tau_{p}$ equals to $\tau_{low}$
Theoretical Result
@@
Under Assumptions 1 and 2, and with appropriate choice of threshold $\tau_{p}$, the sequence of reward models $\{R_{\theta}^{(t)}\}$ converges to a fixed point $R_{\theta}^{\ast}$ with improved accuracy, i.e.,
$\lim_{t\to\infty}\text{Acc}^{(t)}=\text{Acc}^{\ast}\geq\text{Acc}^{(t)}\geq\text{Acc}^{(0)}>0.5.$ (9)
Proof
We provide a proof by induction.
Base Case ( $t=0$ ): By Assumption 1, $\text{Acc}^{(0)}>0.5$.
Inductive Step: Assume that at iteration $t$, $\text{Acc}^{(t)}>0.5$. The filtered dataset $\mathcal{D}_{\text{filtered}}^{(t)}$ consists of examples where the model’s predicted probabilities are confident, i.e., $|p^{(t)}(Q,A)-0.5|\geq\delta_{p}$. From Assumption 2, the accuracy on $\mathcal{D}_{\text{filtered}}^{(t)}$ is at least $\alpha>0.5$.
Retraining the model on $\mathcal{D}_{\text{filtered}}^{(t)}$ leads to an updated model $R_{\theta}^{(t+1)}$ with improved accuracy due to the following reasons: 1. Risk Minimization: Training minimizes the empirical risk on $\mathcal{D}_{\text{filtered}}^{(t)}$, leading to better performance on data similar to $\mathcal{D}_{\text{filtered}}^{(t)}$. 2. Data Distribution Shift: Since $\mathcal{D}_{\text{filtered}}^{(t)}$ is a subset where the model is confident and likely correct, retraining on this data reinforces correct predictions.
Therefore, the expected accuracy satisfies $\text{Acc}^{(t+1)}\geq\text{Acc}^{(t)}$.
Convergence: As $\text{Acc}^{(t)}$ is bounded above by 1 and forms a non-decreasing sequence, it converges to $\text{Acc}^{\ast}\leq 1$. Thus,
$\lim_{t\to\infty}\text{Acc}^{(t)}=\text{Acc}^{\ast}\geq\text{Acc}^{(0)}>0.5.$ (10)
$\blacksquare$
The key to convergence is the selection of high-confidence data that is more likely to be correctly labeled. By ensuring $\alpha>0.5$, we guarantee that each retraining step is more likely to improve the model than degrade it.
@@ A.2 Convergence Properties of PPO with a Learned Reward Model
We now analyze the convergence properties of PPO when using the learned reward model $R_{\theta}^{\ast}$ obtained from self-training. PPO is a policy gradient method that seeks to maximize the expected cumulative reward. The policy $\pi_{\phi}$ is updated to maximize:
$J(\phi)=\mathbb{E}_{Q,A\sim\pi_{\phi}}[R_{\theta}^{\ast}(Q,A)].$ (11)
Assumption 3 (Lipschitz Continuity of Reward Model).
The learned reward model $R_{\theta}^{\ast}(Q,A)$ is Lipschitz continuous with respect to $A$, i.e., there exists $L_{R}>0$ such that for all $A_{1},A_{2}$,
$|R_{\theta}^{\ast}(Q,A_{1})-R_{\theta}^{\ast}(Q,A_{2})|\leq L_{R}\|A_{1}-A_{2}\|_{1}.$ (12)
Assumption 4 (Bounded Policy Updates).
The policy updates satisfy $\|\phi^{(t+1)}-\phi^{(t)}\|_{2}\leq\delta_{\phi}$ for some $\delta_{\phi}>0$.
Theorem 2 (Convergence of PPO with Learned Reward Model).
Under Assumptions 3 and 4, and given that the true reward function $R^{*}(Q,A)$ is approximated by $R_{\theta}^{\ast}(Q,A)$ with bounded error $\epsilon_{r}$:
$|R_{\theta}^{\ast}(Q,A)-R^{*}(Q,A)|\leq\epsilon_{r},$ (13)
the PPO algorithm converges to a policy $\pi_{\phi}^{\ast}$ that is within $\mathcal{O}(\epsilon_{r})$ of the optimal policy $\pi^{*}$ with respect to $R^{*}(Q,A)$.
Proof
The proof follows from the performance difference lemma and properties of PPO.
Performance Difference Lemma(Kakade & Langford, 2002):
The difference in expected rewards between the learned policy $\pi_{\phi}$ and the optimal policy $\pi^{*}$ under the true reward function $R^{*}$ is:
$J^{*}(\pi^{*})-J^{*}(\pi_{\phi})=\frac{1}{1-\gamma}\mathbb{E}_{Q\sim\mu}\left[\mathbb{E}_{A\sim\pi^{*}}\left[A_{\pi^{*}}^{\pi_{\phi}}(Q,A)\right]\right],$ (14)
where $A_{\pi^{*}}^{\pi_{\phi}}(Q,A)$ is the advantage function.
Since $R_{\theta}^{\ast}$ approximates $R^{*}$ with error $\epsilon_{r}$, the advantage estimates used in PPO are off by at most $\epsilon_{r}$.
Impact on Policy Gradient: The policy gradient used in PPO is:
$\nabla_{\phi}J(\phi)=\mathbb{E}_{Q,A\sim\pi_{\phi}}[\nabla_{\phi}\log\pi_{\phi}(A\mid Q)A^{\text{R}}(Q,A)],$ (15)
where $A^{\text{R}}(Q,A)$ is the advantage function computed using $R_{\theta}^{\ast}$.
Due to the bounded reward error $\epsilon_{r}$, the gradient estimation error is also bounded:
$\|\nabla_{\phi}J^{*}(\phi)-\nabla_{\phi}J(\phi)\|_{2}\leq C\epsilon_{r},$ (16)
where $C$ is a constant depending on the policy and reward model.
Convergence Analysis:
Under Assumptions 3 and 4, standard results from stochastic gradient descent convergence apply (Bottou et al., 2018). The policy updates converge to a stationary point of $J(\phi)$, and the error in the reward model introduces an $\mathcal{O}(\epsilon_{r})$ bias.
Therefore, the final policy $\pi_{\phi}^{\ast}$ satisfies:
$|J^{*}(\pi^{*})-J^{*}(\pi_{\phi}^{\ast})|\leq K\epsilon_{r},$ (17)
for some constant $K$.
$\blacksquare$
The convergence to a near-optimal policy depends on the accuracy of the learned reward model. As $\epsilon_{r}\to 0$, the learned policy approaches the optimal policy under $R^{*}$.
Combining Theorems 1 and 2, we conclude that: (1) The reward model improves over iterations, reducing the reward estimation error $\epsilon_{r}$. (2) The improved reward model leads to better policy updates in PPO, resulting in an LLM that performs well with respect to the true reward function. (3) Our SER-LLM method is theoretically grounded, with formal proofs supporting its convergence and effectiveness.
Appendix B Algorithm